# A Path to Universal Neural Cellular Automata

### v1

@@ Abstract.
Cellular automata have long been celebrated for their ability to generate complex behaviors from simple, local rules, with well-known discrete models like Conway’s Game of Life proven capable of universal computation. Recent advancements have extended cellular automata into continuous domains, raising the question of whether these systems retain the capacity for universal computation. In parallel, neural cellular automata have emerged as a powerful paradigm where rules are learned via gradient descent rather than manually designed. This work explores the potential of neural cellular automata to develop a continuous Universal Cellular Automaton through training by gradient descent. We introduce a cellular automaton model, objective functions and training strategies to guide neural cellular automata toward universal computation in a continuous setting. Our experiments demonstrate the successful training of fundamental computational primitives — such as matrix multiplication and transposition — culminating in the emulation of a neural network solving the MNIST digit classification task directly within the cellular automata state. These results represent a foundational step toward realizing analog general-purpose computers, with implications for understanding universal computation in continuous dynamics and advancing the automated discovery of complex cellular automata behaviors via machine learning.
@@
journalyear: 2025copyright: ccconference: Genetic and Evolutionary Computation Conference; July 14–18, 2025; Malaga, Spainbooktitle: Genetic and Evolutionary Computation Conference (GECCO ’25 Companion), July 14–18, 2025, Malaga, Spaindoi: 10.1145/3712255.3734310isbn: 979-8-4007-1464-1/2025/07
@@
 [^1] Web version of this article available at https://gabrielbena.github.io/blog/2025/bena2025unca/
@@ 1. Introduction
Cellular Automata (CA) represent a fascinating class of computational models that have captivated researchers across disciplines for their ability to produce complex behaviors from simple rules (Flake, 1998). At their core, CA typically consist of a grid of cells, each in one of a finite number of states, which evolve over discrete time steps based on a fixed deterministic rule. This rule, applied uniformly to all cells, governs the state transition of each cell based solely on its current state and those of its neighbors. Despite their simplicity, CA have become a cornerstone in studying emergence and complexity (Wolfram, 1984). Mathematical proofs have established that well-known CA, such as Conway’s Game of Life, Elementary Cellular Automata, and Wireworld, are capable of universal computation, underscoring their remarkable power and expressiveness (Department of Computation and Neural Systems, Caltech, Mail Stop 136-93, Pasadena, California 91125, USA and Cook, 2004; Rendell, 2016). Beyond these formal demonstrations, researchers have constructed fully functional Turing machines within these CA, albeit through arduous efforts requiring meticulous design and substantial time investment.
@@
In parallel with these developments, Neural Cellular Automata (NCA) have emerged as a compelling paradigm that blends the local, decentralized dynamics of classical CA with the representational capacity and trainability of neural networks (Mordvintsev et al., 2020). Unlike traditional CA, where the update rule is explicitly handcrafted, NCA leverage differentiable architectures wherein the rule is parameterized by a neural network and optimized end-to-end via gradient descent. This makes it possible to learn complex behaviors and dynamics from data, bypassing the need for manual rule design. Recent work has demonstrated that NCA can be trained to perform a wide range of tasks: from self-organizing into complex morphologies (Mordvintsev et al., 2020), to solving algorithmic reasoning tasks such as instances of the 1D-ARC challenge (Faldor and Cully, 2025), exhibiting emergent collective behaviors (Randazzo et al., 2020), and growing artificial neural networks (Najarro et al., 2022). These results highlight the versatility of NCA as a model of computation and pattern formation. Given the Turing completeness of classical CA, NCA offer an exciting new lens through which to explore the space of rules — not by manually engineering them, but by searching for them through optimization. In essence, NCA turn rule discovery into a machine learning problem. This shift is significant: the traditionally arduous task of hand-crafting rule sets that give rise to desired behaviors is now offloaded to the learning algorithm itself. Backpropagation through time, combined with the differentiable nature of NCA, allows for flexible, automated tuning of highly non-trivial dynamics.
@@
In this work, we explore the potential of the Neural Cellular Automata paradigm to pioneer the development of a continuous Universal Cellular Automata (Weisstein, [n. d.]), with the ambitious goal of inducing a universal Turing machine (Turing, 1937) to emerge within a continuous CA through training via gradient descent. The development of Universal Neural Cellular Automata has implications beyond academic curiosity or creating interesting simulations. It touches on fundamental questions about the potential for continuous dynamic systems to exhibit universal computation, or the possibility to create a universal analog computer. This paper establishes the foundational framework and presents promising initial steps toward realizing this grand vision. First, we propose a novel framework that disentangles the concepts of “hardware” and “state” within NCA. In this abstraction, CA rules serve as the “physics” dictating state transitions across space and time, akin to the fundamental laws governing computation. The CA state, in turn, acts as the dynamic physical substrate — comparable to electrical charges in a computer or neurochemical patterns in a brain—while the “hardware” represents an immutable scaffold, fixed in its spatial configuration throughout a simulation. This hardware can be leveraged by the physics (i.e., the CA rules) to guide computation but remains unalterable during runtime, providing a stable backbone for emergent behaviors. Second, we introduce preliminary objective functions and training setups designed to steer NCA toward universal computation in a continuous domain. Third, we conduct experiments demonstrating the training of essential computational building blocks — such as matrix multiplication, dot-product and transposition — within the NCA framework. Finally, we showcase the practical utility of these building blocks by emulating a neural network directly within the mutable CA state, successfully solving the MNIST digit classification task. These results mark a critical first step, illustrating how NCA can harness gradient descent to sculpt continuous CA into powerful, general-purpose computational systems.
@@ 2. Related Work
Discrete CA have long been a cornerstone for studying universal computation due to their ability to generate complex behaviors from simple, local rules. A number of discrete CA, such as Conway’s Game of Life (Gardner, 1970; Rendell, 2016), Rule 110 of Elementary Cellular Automata (Wolfram, 2002; Department of Computation and Neural Systems, Caltech, Mail Stop 136-93, Pasadena, California 91125, USA and Cook, 2004), Langton’s Ant (Gajardo et al., 2002), and Wireworld (ADA University et al., 2018) have all been proven Turing-complete. These works rely on discrete states and labor-intensive, hand-crafted designs. To mitigate this, evolutionary algorithms have been employed to automate the discovery of CA rules or patterns with specific properties (Mitchell et al., 2000; Sapin et al., 2003), reducing human effort though still targeting discrete systems and predefined goals rather than general-purpose computation.
The shift to continuous CA aims to connect discrete models with real-world analog systems, prompting inquiries into their computational potential. Models like Lenia (Chan, 2019) and SmoothLife (Rafler, 2011) introduce smooth state transitions, yielding lifelike emergent patterns, yet their capacity for universal computation remains unproven. Recent efforts have applied evolutionary search to Lenia (Reinke et al., 2020; len, 2024) to automatically discover and optimize patterns, though these pursuits prioritize specific behaviors over general-purpose computation. Similarly, gradient descent has been used on Lenia patterns and rules to discover and optimize patterns (Hamon et al., 2024), though again not targeting general-purpose computation.
@@ 3.3.1. Monolithic hardware
Figure 3. Monolithic hardware configurations for 3 different sub-tasks. We plot a PCA projection of the hidden channels to be able to display them as RGB. Colors thus does not have direct functional relevance.
@@ 3.3.2. Modular hardware
Figure 4. Modular hardware configurations for 3 different sub-tasks, and different matrix sizes and placement
@@ 4.2. MNIST Classifier emulation
Figure 5. NCAs as a neural network emulator.
Inputs show a batch of flattened MNIST images, alongside the weight matrix of a pre-trained single-layer linear classifier. We decompose this matrix multiplication into sub-blocks, that can be directly emulated in parallel by the NCA. Results are fetched from NCA states and aggregated back into logits, of size batch x 10 (figure shows the first 32 outputs only for readability). We compute accuracy by taking the logits argmax per batch and comparing with labels.
@@ 4.3. Future directions: task composition and neural compiler
This composite task chaining also highlights the crucial role of stability in achieving composable computations. When outputs produced within the NCA’s computational state serve as inputs for subsequent operations, establishing stable representations becomes essential. In parallel to biological systems, where homeostasis maintains internal equilibrium despite external changes, NCAs require a form of computational homeostasis to maintain reliable state representations between operations. Similarly, analog computing systems may require material homeostasis to successfully implement extended chains of tasks and computations without degradation of information.
We propose that this sequential operation model suggests a compelling dual-timestep approach to neural compilation: At the neuronal timestep, the NCA’s mutable state evolves according to its update rules, creating the fundamental dynamics of computation. At the compiler timestep, hardware parameters are reconfigured to provide task abstractions and high-level procedural steps. This separation of concerns, where fast neuronal dynamics handle computation while slower hardware changes control program flow, mirrors classical computer architecture but within a continuous, differentiable substrate. As this approach matures, it could enable direct compilation of algorithms into neural cellular automata, combining the flexibility of neural networks with the precision of programmatic execution. This would possibly be facilitated by a better task-abstraction and hardware generation that we detail in the next session.
@@ 5. Conclusion
The exploration of universal computation within cellular automata has historically been confined to discrete systems, where models like Conway’s Game of Life and Elementary Cellular Automata have demonstrated the remarkable ability to emulate Universal Turing Machines. However, extending this capability to continuous cellular automata presents significant challenges, primarily due to the absence of discrete states and the inherent instability of smooth, analog dynamics. In this work, we have taken pragmatic first steps toward overcoming these hurdles by leveraging NCA as a substrate for developing universal computation in a continuous domain. By employing gradient descent to train NCA rules, we have demonstrated a pathway to sculpt complex computational behaviors without the need for manual rule design, shifting the burden of discovery from human ingenuity to machine learning.
Our results illustrate that NCA can successfully encode fundamental computational primitives, such as matrix multiplication and inversion, and even emulate a neural network capable of solving the MNIST digit classification task directly within its state. These findings suggest that NCAs can serve as a bridge between traditional computing architectures and self-organizing systems, offering a novel computational paradigm that aligns closely with analog hardware systems. This linkage is particularly promising for designing efficient computational frameworks for AI models operating under constrained resources, where energy efficiency and robustness are paramount. Rather than training entirely new rules for each task, our approach hints at the possibility of discovering optimal hardware configurations that exploit the fixed physical laws governing these substrates, enabling meaningful computations with minimal overhead.

### v2

@@ Abstract.
Cellular automata have long been celebrated for their ability to generate complex behaviours from simple, local rules, with well-known discrete models like Conway’s Game of Life proven capable of universal computation. Recent advancements have extended cellular automata into continuous domains, raising the question of whether these systems retain the capacity for universal computation. In parallel, neural cellular automata have emerged as a powerful paradigm where rules are learned via gradient descent rather than manually designed. This work explores the potential of neural cellular automata to develop a continuous Universal Cellular Automaton through training by gradient descent. We introduce a cellular automaton model, objective functions and training strategies to guide neural cellular automata toward universal computation in a continuous setting. Our experiments demonstrate the successful training of fundamental computational primitives — such as matrix multiplication and transposition — culminating in the emulation of a neural network solving the MNIST digit classification task directly within the cellular automata state. These results represent a foundational step toward realizing analog general-purpose computers, with implications for understanding universal computation in continuous dynamics and advancing the automated discovery of complex cellular automata behaviours via machine learning.footnotetext: Web version available at https://gabrielbena.github.io/blog/2025/bena2025unca/
@@
copyright: none
@@
Contributions
Gabriel Béna and Maxence Faldor contributed equally to this work. Maxence Faldor initiated the project, conceived its central ideas, including the concept of state and immutable state for hardware representation. He also developed the initial software library and experimental design. Gabriel Béna provided crucial feedback on these foundational ideas, developed the tasks setup, led the development of the compiler and the Graph Neural Network (GNN) components, conducted the extensive experiments, and was the primary author of the project website. Dan Goodman and Antoine Cully supervised the project, offering guidance, insightful discussions, and contributing to brainstorming sessions.
@@ 1. Introduction
Cellular Automata (CA) represent a fascinating class of computational models that have captivated researchers across disciplines for their ability to produce complex behaviours from simple rules (Flake, 1998). At their core, CA typically consist of a grid of cells, each in one of a finite number of states, which evolve over discrete time steps based on a fixed deterministic rule. This rule, applied uniformly to all cells, governs the state transition of each cell based solely on its current state and those of its neighbors. Despite their simplicity, CA have become a cornerstone in studying emergence and complexity (Wolfram, 1984). Mathematical proofs have established that well-known CA, such as Conway’s Game of Life, Elementary Cellular Automata, and Wireworld, are capable of universal computation, underscoring their remarkable power and expressiveness (Department of Computation and Neural Systems, Caltech, Mail Stop 136-93, Pasadena, California 91125, USA and Cook, 2004; Rendell, 2016). Beyond these formal demonstrations, researchers have constructed fully functional Turing machines within these CA, albeit through arduous efforts requiring meticulous design and substantial time investment.
@@
In parallel with these developments, Neural Cellular Automata (NCA) have emerged as a compelling paradigm that blends the local, decentralized dynamics of classical CA with the representational capacity and trainability of neural networks (Mordvintsev et al., 2020). Unlike traditional CA, where the update rule is explicitly handcrafted, NCA leverage differentiable architectures wherein the rule is parameterized by a neural network and optimized end-to-end via gradient descent. This makes it possible to learn complex behaviours and dynamics from data, bypassing the need for manual rule design. Recent work has demonstrated that NCA can be trained to perform a wide range of tasks: from self-organizing into complex morphologies (Mordvintsev et al., 2020), to solving algorithmic reasoning tasks such as instances of the 1D-ARC challenge (Faldor and Cully, 2025), exhibiting emergent collective behaviours (Randazzo et al., 2020), and growing artificial neural networks (Najarro et al., 2022). These results highlight the versatility of NCA as a model of computation and pattern formation. Given the Turing completeness of classical CA, NCA offer an exciting new lens through which to explore the space of rules — not by manually engineering them, but by searching for them through optimization. In essence, NCA turn rule discovery into a machine learning problem. This shift is significant: the traditionally arduous task of hand-crafting rule sets that give rise to desired behaviours is now offloaded to the learning algorithm itself. Backpropagation through time, combined with the differentiable nature of NCA, allows for flexible, automated tuning of highly non-trivial dynamics.
@@
In this work, we explore the potential of the Neural Cellular Automata paradigm to pioneer the development of a continuous Universal Cellular Automata (Weisstein, [n. d.]), with the ambitious goal of inducing a universal Turing machine (Turing, 1937) to emerge within a continuous CA through training via gradient descent. The development of Universal Neural Cellular Automata has implications beyond academic curiosity or creating interesting simulations. It touches on fundamental questions about the potential for continuous dynamic systems to exhibit universal computation, or the possibility to create a universal analog computer. This paper establishes the foundational framework and presents promising initial steps toward realizing this grand vision. First, we propose a novel framework that disentangles the concepts of “hardware” and “state” within NCA. In this abstraction, CA rules serve as the “physics” dictating state transitions across space and time, akin to the fundamental laws governing computation. The CA state, in turn, acts as the dynamic physical substrate — comparable to electrical charges in a computer or neurochemical patterns in a brain—while the “hardware” represents an immutable scaffold, fixed in its spatial configuration throughout a simulation. This hardware can be leveraged by the physics (i.e., the CA rules) to guide computation but remains unalterable during runtime, providing a stable backbone for emergent behaviours. Second, we introduce preliminary objective functions and training setups designed to steer NCA toward universal computation in a continuous domain. Third, we conduct experiments demonstrating the training of essential computational building blocks — such as matrix multiplication, dot-product and transposition — within the NCA framework. Finally, we showcase the practical utility of these building blocks by emulating a neural network directly within the mutable CA state, successfully solving the MNIST digit classification task. These results mark a critical first step, illustrating how NCA can harness gradient descent to sculpt continuous CA into powerful, general-purpose computational systems.
@@ 2. Related Work
Discrete CA have long been a cornerstone for studying universal computation due to their ability to generate complex behaviours from simple, local rules. A number of discrete CA, such as Conway’s Game of Life (Gardner, 1970; Rendell, 2016), Rule 110 of Elementary Cellular Automata (Wolfram, 2002; Department of Computation and Neural Systems, Caltech, Mail Stop 136-93, Pasadena, California 91125, USA and Cook, 2004), Langton’s Ant (Gajardo et al., 2002), and Wireworld (ADA University et al., 2018) have all been proven Turing-complete. These works rely on discrete states and labor-intensive, hand-crafted designs. To mitigate this, evolutionary algorithms have been employed to automate the discovery of CA rules or patterns with specific properties (Mitchell et al., 2000; Sapin et al., 2003), reducing human effort though still targeting discrete systems and predefined goals rather than general-purpose computation.
The shift to continuous CA aims to connect discrete models with real-world analog systems, prompting inquiries into their computational potential. Models like Lenia (Chan, 2019) and SmoothLife (Rafler, 2011) introduce smooth state transitions, yielding lifelike emergent patterns, yet their capacity for universal computation remains unproven. Recent efforts have applied evolutionary search to Lenia (Reinke et al., 2020; len, 2024) to automatically discover and optimize patterns, though these pursuits prioritize specific behaviours over general-purpose computation. Similarly, gradient descent has been used on Lenia patterns and rules to discover and optimize patterns (Hamon et al., 2024), though again not targeting general-purpose computation.
@@ 3.3.1. Monolithic hardware
Figure 3. Monolithic hardware configurations for 3 different sub-tasks. We plot a PCA projection of the hidden channels to be able to display them as RGB. Colors thus does not have direct functional relevance. Overlayed are schematics of the (fixed) input-output transformations that each hardware was optimized on.
@@ 3.3.2. Modular hardware
Figure 4. Modular hardware configurations for 3 different sub-tasks, and (now varying) different matrix sizes and placement.
@@ 4.2. MNIST Classifier emulation
Figure 5. NCA emulates a neural network. Inputs show a batch of flattened MNIST images, alongside the weight matrix of a pre-trained single-layer linear classifier. We decompose this matrix multiplication into sub-blocks, that can be directly emulated in parallel by the NCA. Results are fetched from NCA states and aggregated back into logits (figure shows the first 32 outputs only for readability). We compute accuracy by taking the logits argmax per batch and comparing with labels.
@@ 4.3. Future directions: task composition and neural compiler
This composite task chaining also highlights the crucial role of stability in achieving composable computations. When outputs produced within the NCA’s computational state serve as inputs for subsequent operations, establishing stable representations becomes essential. In parallel to biological systems, where homeostasis maintains internal equilibrium despite external changes, NCAs require a form of computational homeostasis to maintain reliable state representations between operations. Similarly, analog computers may require homeostasis to successfully implement extended chains of tasks and computations without degradation of information.
We propose that this sequential operation model suggests a compelling dual-timestep approach to neural compilation: At the neuronal timestep, the NCA’s mutable state evolves according to its update rules, creating the fundamental dynamics of computation. At the compiler timestep, hardware parameters are reconfigured to provide task abstractions and high-level procedural steps. This separation of concerns, where fast neuronal dynamics handle computation while slower hardware changes control program flow, mirrors classical computer architecture but within a continuous, differentiable substrate. As this approach matures, it could enable direct compilation of algorithms into neural cellular automata, combining the flexibility of neural networks with programmatic execution. This would possibly be facilitated by a better task-abstraction and hardware generation that we detail in the next section.
@@ 5. Conclusion
The exploration of universal computation within cellular automata has historically been confined to discrete systems, where models like Conway’s Game of Life and Elementary Cellular Automata have demonstrated the remarkable ability to emulate Universal Turing Machines. However, extending this capability to continuous cellular automata presents significant challenges, primarily due to the absence of discrete states and the inherent instability of smooth, analog dynamics. In this work, we have taken pragmatic first steps toward overcoming these hurdles by leveraging NCA as a substrate for developing universal computation in a continuous domain. By employing gradient descent to train NCA rules, we have demonstrated a pathway to sculpt complex computational behaviours without the need for manual rule design, shifting the burden of discovery from human ingenuity to machine learning.
Our results illustrate that NCA can successfully encode fundamental computational primitives, such as matrix multiplication and inversion, and even emulate a neural network capable of solving the MNIST digit classification task directly within its state. These findings suggest that NCAs can serve as a bridge between traditional computing architectures and self-organizing systems, offering a novel computational paradigm that aligns closely with analog systems. This linkage is particularly promising for designing efficient computational frameworks for AI models operating, where energy efficiency and robustness are paramount. Rather than training entirely new rules for each task, our approach hints at the possibility of discovering optimal hardware configurations that exploit the fixed physical laws governing these substrates, enabling meaningful computations with minimal overhead.